OpenStudy (anonymous):

For the given cost function C(x)=28900+400x+x^2 find: -the production level that will minimize average cost -the minimal average cost

OpenStudy (anonymous):

set up the 1st derivative of the function and set it equal to zero, then solve for x... C(x)=28900+400x+x^2 C'(x) = 400 + 2x C'(x) = 0 400 + 2x = 0 2x = -400 x = -200 the cost will minimum when x = -200

OpenStudy (anonymous):

thank you